# Polar Mounds on Strangeon Stars: the Neutrino Emission from Ultraluminous X-ray Pulsars

### v1

@@ 2.2 Calculation of the polar mound height
A key question in ULXPs is the region capable of effectively radiating neutrinos. A. A. Mushtukov et al. (2018) proposed that strong advection at extremely high mass accretion rates leads to the formation of a zone in the central region of the accretion column, where the released energy is carried downward and emitted as neutrinos. In their study, the height of the neutrino emission region is determined by photon advection. However, their model did not consider the thermal mound properties, nor did it address how the structure and EOS of the pulsar could significantly influence the thermal mound characteristics and the efficiency of neutrino emission.
@@
In this work, we investigate the thermal mound properties based on SS model (R.-X. Xu, 2003), taking into account both the Coulomb barrier (R. X. Xu & G. J. Qiao, 1999; J. Hu & R. X. Xu, 2002; R.-X. Xu, 2014) and the strangeness barrier (R.-X. Xu, 2014; Z. Li et al., 2015; W. Wang et al., 2017; W.-Y. Wang et al., 2018). The inflowing matter penetrates the Coulomb barrier and follows the mass conservation equation,
$\dot{M}(R_{M})=\tau\frac{{\rm d}N_{n}}{{\rm d}t}m_{p}\,$, (7)
@@
where $\dot{M}(R_{M})$ is the mass accretion rate at the magnetospheric radius, $m_{p}$ is the proton mass, and $\tau$ is the penetration probability, given by
$\tau=\exp\left[-\left(\frac{E_{\rm G}}{E}\right)^{1/2}\right]\,$, (8)
@@
with $E_{\rm G}=0.49\,\rm MeV$ as the Gamow energy (D. H. Perkins, 2003), and $E=k_{B}T_{\rm C}$, where $k_{B}$ is the Boltzmann constant and $T_{\rm C}$ is the characteristic temperature, which depends on the mass accretion rate (see Fig. 2). The penetration time of the nucleons is approximately $t=0.01c/(1/n)^{1/3}\approx 1\times 10^{17}\,\rm s$, and the number of nucleons attempting to penetrate is $N_{n}=\lambda nSh$, where $n$ is the number density, $S=\pi R_{\rm S}^{2}\approx 10^{10}\,\rm cm^{2}$ is the cross-sectional area of the accretion column base, the constant $\lambda=0.01$, and $h\approx(1/n)^{1/3}$. Our results indicate that the Coulomb barrier in strangeon stars does not effectively prevent the inflowing matter from penetrating the star.
@@
Furthermore, the falling matter may significantly penetrate the strangeness barrier, most of the falling matter would be bounced back along the magnetic field lines because nonstrange matter cannot become part of strange matter unless it is converted to strangeons via weak interaction (R.-X. Xu, 2014; W. Wang et al., 2017). This process is key to the formation of the thermal mound.
To estimate the thermal mound height, we first calculate the number density of nonstrange normal matter at the base of the polar cone. Assuming the weak interaction timescale is $t^{*}_{\rm weak}=10^{-10}\,\rm s$, the probability of ions changing flavor to strangeon is approximately $\eta\sim 10^{-13}$, and the speed of ion is $v_{i}\sim 0.01c$. The number density at the base of the polar mound is related to the mass accretion rate via
@@
which is related to EOS. We consider contributions from gas pressure, radiation pressure, and degenerate pressure, as described in B. Paczynski (1983). Using Eq. (10) in conjunction with EOS, we can obtain the height of the thermal mound.
The upper panel of Fig. 3 shows the mass density of the thermal mound as a function of thermal mound height for different weak interaction timescales, with a fixed mass accretion rate $\dot{M}=\dot{M}_{\rm Edd}$. The thermal mound height $H$ ranges from $0.7$ to $0.95,\rm km$. The lower panel of Fig. 3 shows the mass density as a function of thermal mound height for different mass accretion rates. As the mass accretion rate increases, the thermal mound height becomes larger.
The density at the base of the thermal mound in SSs is approximately $10^{11}\,\rm g\,cm^{-3}$, while the surface density of strangeon matter exceeds $10^{14}\,\rm g\,cm^{-3}$, resulting in a discontinuity in mass density at the surface. This discontinuity reduces the number of particles in the polar mound region of an SS, lowering its effective heat capacity. With thermal energy supplied by gravitational potential energy, the reduced heat capacity leads to a sharply elevated temperature, which, together with the high density, promotes efficient neutrino production. In contrast, the surface density of an NS is continuous and generally lower than $10^{11}\,\rm g\,cm^{-3}$ (D. Mukherjee & D. Bhattacharya, 2012; D. Mukherjee, 2017), and the thermal mound height in NSs is significantly smaller than in SSs (D. Mukherjee et al., 2013a, b). In an NS, the same thermal energy is distributed over more matter, resulting in a higher heat capacity and a much lower polar mound temperature, rendering neutrino emission inefficient. These contrasts highlight the distinctive thermal and neutrino emission properties of SSs relative to NSs.
Figure 3: Mass density of the thermal mound as a function of thermal mound height for different weak interaction timescales with a fixed mass accretion rate $\dot{M}=\dot{M}_{\rm Edd}$ (upper panel), and for different mass accretion rates with a fixed weak interaction timescale $t^{*}_{\rm weak}=10^{-10}\,\rm s$ (lower panel).
@@ 3 Heat transport from strangeon star surface
In this section, we discuss the heat energy transport from SS surface. According to our results, the temperature at the base of the polar mound exceeds $10^{9}\,\rm K$, and this thermal energy can be transported along the SS surface to other regions. The heat flux is related to the temperature gradient through Fourier’s law, ${\bf F}=-\kappa\nabla T$, where the thermal conductivity $\kappa$ is primarily determined by electron-phonon interactions. The typical value is $1.2\times 10^{20}\,\rm erg\,K^{-1}\,cm^{-1}\,s^{-1}$ (E. Flowers & N. Itoh, 1976; M. Yu & R. X. Xu, 2009).
In this work, we adopt a one-dimensional approximation for the temperature gradient, $\nabla T\simeq(T_{\rm C}-T_{p})/R_{\rm s}$, where $T_{\rm C}$ is the base temperature of the polar mound (see Fig. 2), $T_{p}\simeq 2-3\,\rm keV$ (M. Middleton et al., 2015; M. Kumar et al., 2025). Assuming that the high-temperature region corresponds to the base of the accretion column, the luminosity associated with heat transport can be expressed as
@@
4 Neutrino flux and and detection
@@
where $L_{\rm ph}$ and $L_{\rm\nu}$ are the neutrino and photon luminosity, respectively. Following the methodology of A. A. Mushtukov et al. (2018); A. Asthana et al. (2023), the neutrino luminosity is given by
$L_{\rm\nu}\approx S\int^{H}_{0}Q_{\rm acc}(H)dH\,$, (13)
where $H$ is the hight of the thermal mound, which is dependent on the weak interaction timescale and mass accretion rate (see Fig. 3). The accretion heating rate defined by variations of the gravitational potential energy is given by
$Q_{\rm acc}(H)=\frac{\dot{M}}{2S}\left[\frac{GM}{(R+H)^{2}}\right]\,.$ (14)
@@
Figure 4: The total (dash-dotted line), the neutrino ($L_{\nu}$) and photon luminosity ($L_{\rm ph}$) as function of the mass accretion rate. The horizontal dash lines are show the observed luminosity for some ULXPs. We set $t=t^{*}_{\rm weak}$.
In Fig. 4, we plot the total, neutrino and photon luminosity as functions of the mass accretion rate. At relatively low mass accretion rates ($10^{20}\,\rm g\,s^{-1}$), photons can escape from the accretion column, and the neutrino luminosity remains negligible, with the total luminosity being dominated by photons. However, at higher mass accretion rates ($10^{21}\,\rm g\,s^{-1}$), a large region forms where photons are trapped, and energy is transported to regions cooled by neutrinos, leading to a total luminosity that exceeds the photon luminosity. Furthermore, as the mass accretion rate increases, most of the energy is converted into neutrinos, causing the photon luminosity growth to slow. The photon luminosity may approach an upper limit of around $10^{41}\,\rm erg\,s^{-1}$.
@@
We neglect neutrino oscillations in the vicinity of the X-ray source due to the significant uncertainty in the geometry of the accretion column and its small extent. The spectra of different neutrino flavors are assumed to be the same, as the kinematics of the annihilation process are identical (see, for example, A. Asthana et al. (2023) for a detailed discussion).
The photon luminosity of the sources, expected neutrino luminosity, distance to the source, and corresponding neutrino energy flux at Earth from ULXPs are summarized in Table 2. It can be seen that, except for Swift J0243.6++6124, all detected ULXPs are extragalactic, leading to significant signal attenuation due to their large distances. We discuss the possibility of neutrino detection from Swift J0243.6++6124, for which the neutrino flux at Earth is expected to be the highest.
The population synthesis studies suggest that $\sim 7-30$ ULXPs in Milky Way (Y. Shao & X.-D. Li, 2015; S.-J. Gao et al., 2022; L. Li et al., 2024), and $\sim 384-629$ ULX candidates in nearby galaxies (H. P. Earnshaw et al., 2019; K. Kovlakas et al., 2020). Using the results of the population synthesis, we estimate the energy spectrum of neutrino background at Earth. In Fig. 5, we show the energy spectrum of neutrino background at Earth in the energy range from $10^{4}\,\rm eV$ to $10^{12}\,\rm eV$. Our results indicate that even for source Swift J0243.6++6124, detection with current neutrino observatories would be challenging.
@@
Figure 5: The energy spectrum of neutrino background at Earth in the energy range from $10^{4}\,\rm eV$ to $10^{12}\,\rm eV$. The background component of the neutrino flux includes contributions from atmospheric neutrinos, nuclear reactors, old supernovae, and geoneutrinos. These data were taken from E. Vitagliano et al. (2020). Two rectangles highlight the regions where we expect the neutrino flux from the closest source, Swift J0243.6++6124, and from the brightest source, NGC 5907 ULX-1.
@@
Table 2: The photon luminosity of the objects $L_{\rm ph}$, expected neutrino luminosity $L_{\nu}$, distance to the source, and corresponding neutrino energy flux $F_{\nu}$ at Earth from ULXPs.

### v2

@@ 2.2 Calculation of the polar mound height
A key question in ULXPs is the region capable of effectively radiating neutrinos. Mushtukov et al. (2018) proposed that strong advection at extremely high mass accretion rates leads to the formation of a zone in the central region of the accretion column, where the released energy is carried downward and emitted as neutrinos. In their study, the height of the neutrino emission region is determined by photon advection. However, their model did not consider the thermal mound properties, nor did it address how the structure and EOS of the pulsar could significantly influence the thermal mound characteristics and the efficiency of neutrino emission.
@@
In this work, we investigate the thermal mound properties based on SS model (Xu, 2003), taking into account both the Coulomb barrier (Xu and Qiao, 1999; Hu and Xu, 2002; Xu, 2014) and the strangeness barrier (Xu, 2014; Li et al., 2015; Wang et al., 2017, 2018). The inflowing matter penetrates the Coulomb barrier and follows the mass conservation equation,
$\dot{M}(R_{M})=\tau_{\rm cou}\frac{{\rm d}N_{n}}{{\rm d}t}m_{p}\,$, (9)
@@
where $\dot{M}(R_{M})$ is the mass accretion rate at the magnetospheric radius, $m_{p}$ is the proton mass, and $\tau_{\rm cou}$ is the penetration probability, given by
$\tau_{\rm cou}=\exp\left[-\left(\frac{E_{\rm G}}{E}\right)^{1/2}\right]\,$, (10)
@@
with $E_{\rm G}=0.49\,\rm MeV$ as the Gamow energy (Perkins, 2003), and $E=k_{B}T_{\rm base}$, where $k_{B}$ is the Boltzmann constant and $T_{\rm base}$ is the base temperature of the accretion column, which depends on the mass accretion rate (see Table 3). The penetration time of the nucleons is approximately $t=0.01c/(1/n)^{1/3}\approx 1\times 10^{17}\,\rm s$. The number of nucleons attempting to penetrate is $N_{n}=\lambda nSh$, where $n$ is the number density, $h\approx(1/n)^{1/3}$. $S=\pi R_{\rm S}^{2}\approx 10^{10}\,\rm cm^{2}$ is the cross-sectional area of the accretion column base. However, the size of the cross-sectional area of the accretion column base remains a topic of ongoing debate (Lamb et al., 1973; Arons and Lea, 1976, 1980; Elsner and Lamb, 1977, 1976, 1984; Michel, 1977b, a, c; Ghosh et al., 1977; Petterson, 1977b, c, a; Becker, 1998; Becker and Wolff, 2007, 2022; West et al., 2017a, b). In the works of Becker and Wolff (2007) and Becker and Wolff (2022), the shape of the X-ray continuum in accretion-powered pulsars has been successfully fitted using the accretion column model. The cross-sectional area of the accretion column in their models is smaller than that in Asthana et al. (2023). Therefore, we set the parameter $\lambda=0.01$ to modify the size of the cross-sectional area. Our results indicate that the Coulomb barrier in strangeon stars does not effectively prevent the inflowing matter from penetrating the star.
@@
Furthermore, the falling matter may significantly penetrate the strangeness barrier, most of the falling matter would be bounced back along the magnetic field lines because nonstrange matter cannot become part of strange matter unless it is converted to strangeons via weak interaction (Xu, 2014; Wang et al., 2017). This process is key to the formation of the thermal mound.
To estimate the thermal mound height, we first calculate the number density of nonstrange ions accumulated at the base of the polar cone. As discussed earlier, the conversion between $u$, $d$, and $s$ quarks is governed by weak interactions. Motivated by the intrinsic weak decay lifetime of strange baryons such as the $\Lambda$ hyperon, $\tau_{\Lambda}\approx 2.6\times 10^{-10}\,\rm s$, we adopt a typical weak interaction timescale of $t^{*}_{\rm weak}=10^{-10}\,\rm s$. Under this condition, the probability for incident ions to undergo flavor conversion into strangeons is estimated to be $\eta\sim 10^{-13}$, assuming an inflow velocity of $v_{i}\sim 0.01c$. The number density at the base of the polar mound is related to the mass accretion rate via
@@
which is related to EOS. We consider contributions from gas pressure, radiation pressure, and degenerate pressure, as described in Paczynski (1983). Using Eq. (12) in conjunction with EOS, we can obtain the height of the thermal mound. We note that magnetic pressure, $P_{\rm mag}=B^{2}/(8\pi)$, although important for confining the column globally, varies negligibly over the small height of the polar mound ($H_{\rm mound}\approx 1\,\rm km$). Its contribution to the pressure gradient is therefore very small and can be safely neglected in estimating the mound height.
The upper panel of Fig. 3 shows the mass density of the thermal mound as a function of thermal mound height for different weak interaction timescales, with a fixed mass accretion rate $\dot{M}=\dot{M}_{\rm Edd}$. The thermal mound height $H_{\rm mound}$ ranges from $0.7$ to $0.95\,\rm km$. The lower panel of Fig. 3 shows the mass density as a function of thermal mound height for different mass accretion rates. As the mass accretion rate increases, the thermal mound height becomes larger.
The density at the base of the thermal mound in SSs is approximately $10^{11}\,\rm g\,cm^{-3}$, while the surface density of strangeon matter exceeds $10^{14}\,\rm g\,cm^{-3}$, resulting in a discontinuity in mass density at the surface. This discontinuity reduces the number of particles in the polar mound region of an SS, lowering its effective heat capacity. With thermal energy supplied by gravitational potential energy, the reduced heat capacity leads to a sharply elevated temperature, which, together with the high density, promotes efficient neutrino production. In contrast, the surface density of an NS is continuous and generally lower than $10^{11}\,\rm g\,cm^{-3}$ (Mukherjee and Bhattacharya, 2012; Mukherjee, 2017), and the thermal mound height in NSs is significantly smaller than in SSs (Mukherjee et al., 2013a, b). In an NS, the same thermal energy is distributed over more matter, resulting in a higher heat capacity and a much lower polar mound temperature, rendering neutrino emission inefficient. These contrasts highlight the distinctive thermal and neutrino emission properties of SSs relative to NSs.
Figure 3: Mass density of the thermal mound as a function of thermal mound height for different weak interaction timescales with a fixed mass accretion rate $\dot{M}=\dot{M}_{\rm Edd}$ (upper panel), and for different mass accretion rates with a fixed weak interaction timescale $t^{*}_{\rm weak}=10^{-10}\,\rm s$ (lower panel). We set magnetic field strength $B=10^{12}\,\rm G$.
@@ 3 Heat transport from strangeon star surface
In this section, we discuss the heat energy transport from SS surface. According to our results, the temperature at the base of the polar mound exceeds $10^{9}\,\rm K$, and this thermal energy can be transported along the SS surface to other regions. The heat flux is related to the temperature gradient through Fourier’s law, ${\bf F}=-\kappa\nabla T$, where the thermal conductivity $\kappa$ is primarily determined by electron-phonon interactions. The typical value is $1.2\times 10^{20}\,\rm erg\,K^{-1}\,cm^{-1}\,s^{-1}$ (Flowers and Itoh, 1976; Yu and Xu, 2009).
In this work, we adopt a one-dimensional approximation for the temperature gradient, $\nabla T\simeq(T_{\rm base}-T_{p})/R_{\rm s}$, where $T_{\rm base}$ is the base temperature of the accretion column (see Table 3), $T_{p}\simeq 2-3\,\rm keV$ (Middleton et al., 2015; Kumar et al., 2025). Assuming that the high-temperature region corresponds to the base of the accretion column, the luminosity associated with heat transport can be expressed as
@@
4 Neutrino flux and detection
@@
where $L_{\rm ph}$ and $L_{\rm\nu}$ are the neutrino and photon luminosity, respectively. Following the methodology of Mushtukov et al. (2018); Asthana et al. (2023), the neutrino luminosity is given by
$L_{\rm\nu}\approx S\int^{H_{\rm mound}}_{0}Q_{\rm acc}(H_{\rm mound})dH_{\rm mound}\,$, (15)
where $H_{\rm mound}$ is the height of the thermal mound, which is dependent on the weak interaction timescale and mass accretion rate (see Fig. 3). The accretion heating rate defined by variations of the gravitational potential energy is given by
$Q_{\rm acc}(H_{\rm mound})=\frac{\dot{M}}{2S}\left[\frac{GM}{(R+H_{\rm mound})^{2}}\right]\,.$ (16)
@@
Figure 4: The neutrino luminosity of our work (solid line), neutrino luminosity of Asthana et al. (2023) (dashed line), and the photon luminosity as functions of the mass accretion rate. The dash-dotted line is total luminosity is the dash-dotted line. The horizontal dash lines are show the observed luminosity for some ULXPs. We calculate the neutrino luminosity with $t=t^{*}_{\rm weak}$ and the magnetic field strength $B=10^{12}\,\rm G$ in our work.
In Fig. 4, we show the total, neutrino, and photon luminosity as functions of the mass accretion rate. We observe that the results are similar in the following scenarios. At relatively low mass accretion rates ($10^{20}\,\rm g\,s^{-1}$), photons can escape from the accretion column, and the neutrino luminosity remains negligible, with the total luminosity being dominated by photons. However, at higher mass accretion rates ($10^{21}\,\rm g\,s^{-1}$), a large region forms where photons are trapped, and energy is transported to regions cooled by neutrinos, leading to a total luminosity that exceeds the photon luminosity. Furthermore, as the mass accretion rate increases, most of the energy is converted into neutrinos, causing the photon luminosity growth to slow. The photon luminosity may approach an upper limit of around $10^{41}\,\rm erg\,s^{-1}$.
For comparison with the NS model, we also compute the neutrino luminosity from the accretion column using Eq. (1) of Asthana et al. (2023) (dashed lines in Fig. 4). At relatively low accretion rates ($10^{20}\,\rm g\,s^{-1}$), the SS model predicts moderately higher neutrino fluxes for Swift J0243.6++6124, RX J0209.6-7427, and SMC X-3, exceeding the results of Asthana et al. (2023) by factors of $\sim 2-5$. In contrast, for ULXPs, the SS model yields neutrino fluxes that are comparable to or lower than those of the NS model, with typical reductions of $\lesssim 2$ and up to a factor of $\sim 5$ for extreme sources such as NGC 5907 ULX-1. This comparison suggests that, within the SS framework, nearby Be X-ray binaries (e.g. Swift J0243.6++6124 and RX J0209.6-7427) are more promising neutrino sources than ULXPs.
@@
We neglect neutrino oscillations in the vicinity of the X-ray source due to the significant uncertainty in the geometry of the accretion column and its small extent. The spectra of different neutrino flavors are assumed to be the same, as the kinematics of the annihilation process are identical (see, for example, Asthana et al. (2023) for a detailed discussion).
The photon luminosity of the sources, expected neutrino luminosity, distance to the source, and corresponding neutrino energy flux at Earth from ULXPs are summarized in Table 2. It is evident that the inferred neutrino luminosity of Swift J0243.6++6124 is significantly higher than that of NGC 5907 ULX-1, primarily due to the large difference in their distances. Swift J0243.6++6124 is located at only $0.007$ Mpc, while NGC 5907 ULX-1 lies at $17.1$ Mpc. Since the observed neutrino flux decreases with the square of the distance, this naturally results in a discrepancy of nearly four orders of magnitude. Except for Swift J0243.6++6124, all detected ULXPs are extragalactic, leading to substantial attenuation of their signals due to their large distances. We discuss the possibility of neutrino detection from Swift J0243.6++6124, for which the neutrino flux at Earth is expected to be the highest. For various ULX Pulsar sources, the proposed model encompasses a broad range of neutron star surface magnetic field strengths, from $10^{11}\,\rm G$ to $10^{14}\,\rm G$ (Fürst and others, 2016; Fuerst and others, 2018, 2023; King et al., 2023; Fürst et al., 2025). In Table 3, we present the effects of the magnetic field on the neutrino luminosity and corresponding neutrino energy flux for NGC 5907 ULX-1. The neutrino luminosity and corresponding neutrino energy flux change with increasing magnetic field strength, and we hope that other sources will exhibit similar properties.
To estimate an upper limit on the neutrino contribution from the broader ULXP population, we adopt a simple scaling approach. Current observational constraints suggest that there are approximately $\sim 7-30$ ULXPs in the Milky Way (Shao and Li, 2015; Gao et al., 2022; Li et al., 2024) and $\sim 384-629$ ULX candidates in nearby galaxies (Earnshaw et al., 2019; Kovlakas et al., 2020). Under the simplifying assumption that these systems share similar intrinsic properties, we scale the neutrino flux of Swift J0243.6++6124 by a factor of $\sim 7-30$ to represent the Galactic contribution, and the flux of NGC 5907 ULX-1 by a factor of $\sim 384-629$ to represent the contribution from nearby galaxies. We emphasize that this procedure does not constitute a population-synthesis calculation, but rather provides an optimistic upper-limit estimate of the total neutrino flux. Recent LHAASO results indicating $\sim 10$ Galactic ULXs are consistent with this estimate within current uncertainties (LHAASO Collaboration, 2024; Peretti et al., 2025; Kaci et al., 2025).
[…]
The predicted neutrino flux spectra at Earth from the two ULXPs (Swift J0243.6++6124 and NGC 5907 ULX-1) are shown in Fig. 5. The focus on these two systems is motivated by their distinct observational advantages. As discussed earlier, the observable neutrino flux decreases with the square of the distance. Swift J0243.6++6124 is by far the closest ULXP and therefore represents the most promising target for potential neutrino detection. Conversely, the neutrino flux is positively correlated with the source luminosity, and NGC 5907 ULX-1 is one of the brightest ULXs known, with an X-ray luminosity reaching $500\,L_{\rm Edd}$. These considerations make the two sources representative of the most favorable detection scenarios among currently confirmed ULXPs.
Even if additional nearby sources similar to Swift J0243.6++6124 exist but are beamed away from Earth, their neutrino emission would not be observable. Therefore, the total neutrino flux from Galactic ULXPs remains low compared with the diffuse background, and no significant detection is expected with current neutrino observatories. This highlights Swift J0243.6++6124 as the most promising individual target for neutrino detection.
@@
Figure 5: The energy spectrum of neutrino background at Earth in the energy range from $10^{4}\,\rm eV$ to $10^{12}\,\rm eV$. The background component of the neutrino flux includes contributions from atmospheric neutrinos, nuclear reactors, old supernovae, and geoneutrinos. These data were taken from Vitagliano et al. (2020). The open solid rectangles show the neutrino fluxes predicted by the SS model in the MeV energy band. For PULXs, Case A corresponds to Swift J0243.6++6124 and RX J0209.6-7427, while Case B includes NGC 5907 ULX-1 and M51 ULX-8. The open dashed rectangles represent the neutrino fluxes of Be X-ray pulsars and ULX pulsars, respectively, with values taken from Asthana et al. (2023).
@@
Table 2: The photon luminosity of the objects $L_{\rm ph}$, expected neutrino luminosity $L_{\nu}$, distance to the source, and corresponding neutrino energy flux $F_{\nu}$ at Earth from ULXPs. We set magnetic field strength $B=10^{12}\,\rm G$.
@@
Table 3: The base temperature of accretion column $T_{\rm base}$, expected neutrino luminosity $L_{\nu}$, distance to the source, and corresponding neutrino energy flux $F_{\nu}$ of the brightest source, NGC 5907 ULX-1 with the differents magnetic fields.
B $T_{\rm base}$ $L_{\nu}$ $d$ $F_{\nu}$
(G) $(10^{9}\,\rm K)$ $(10^{39}\,\rm erg\,s^{-1})$ $(\rm Mpc)$ $(\rm 10^{-6}\,cm^{-2}\,s^{-1}\,MeV^{-1})$
$10^{12}$ 3 64 $17.1$ 1.14
$10^{13}$ 10 139 $17.1$ 2.64
$10^{14}$ 32 216 $17.1$ 4.12